Nelson Corp. is considering the purchase of a new piece of equipment. The cost savings from the equipment would result in an annual increase in cash flow of $159,400. The equipment will have an initial cost of $475,000 and have a 5 year life. If the salvage value of the equipment is estimated to be $115,000, what is the accounting rate of return? Ignore income taxes.

33.56%.

35.23%.

18.40%.

15.90%
Answer #1

Net income = 159400-(475000-115000/5) = 87400

Accounting rate of return = Net income/Initial investment = 87400/475000 = 18.4%

So answer is c) 18.4%